statistics

A major hurricane is a hurricane with wund speeds of 111 miles per hour or greater. During the last century, the mean number of major hurricanes to strike a certaincountry's mainland per year was about .43. Find the probability that in a given year A exactly one major hurricane will strike the mainland, B at most one hurricanewill strike the mainland, C mor than one major hurricane will strike mainland
Answer #1
Given this

m = 0.65

a. P[1] = e^-0.65*0.65 = .3393
b. P[=1] e^-0.65(1+0.65) = .8614
c. P[>1] = 1 - .8614 = .1386
answered by: Gintare
Answer #2
Given this

m = 0.65

a. P[1] = e^-0.65*0.65 = .3393
b. P[=1] e^-0.65(1+0.65) = .8614
c. P[>1] = 1 - .8614 = .1386
answered by: kezia